OpenStudy (lili.gallegos):

Use implicit differentiation to find an equation of the tangent line to the curve. x^5y^5=32, tangent at (2,1)

OpenStudy (freckles):

know product rule, power rule, constant rule, and chain rule?

OpenStudy (freckles):

or if you want to make this really easy take the 5th root of both sides

rvc (rvc):

if you know all the three rules mentioned by freckles then you can easily solve it :)

rvc (rvc):

if you take the fifth root the equation would be : xy=2

rvc (rvc):

slope : \(\Large \frac{dy}{dx}\)

OpenStudy (lili.gallegos):

Can one of you walk me through this ? Step by step ?

rvc (rvc):

equation of tangent : \[\rm y-y_{1}=\frac{ dy }{ dx }(x-x_{1})\]

OpenStudy (lili.gallegos):

What about the derivative ?

OpenStudy (freckles):

you need to apply the rules (or at least some) to differentiate xy=2

OpenStudy (freckles):

xy is a product so you would use the product rule to differentiate xy

OpenStudy (freckles):

2 is a constant the constant rule says derivative of a constant is 0

OpenStudy (freckles):

(xy)'=0 can you apply the product rule to find (xy)'

rvc (rvc):

you can go through this : Basic Derivatives http://openstudy.com/study#/updates/5550950fe4b061b8579d7c1d

OpenStudy (lili.gallegos):

What about the derivative ?

OpenStudy (freckles):

to differentiate is to find the derivative

OpenStudy (freckles):

(xy)' means find the derivative of the product of x and y

OpenStudy (freckles):

do you know how to find the derivative of the product of x and y

OpenStudy (freckles):

do you know the product rule?

OpenStudy (freckles):

actually if you don't know the product rule @rvc has posted for you in that link

OpenStudy (lili.gallegos):

Yes I know the product rule ..

OpenStudy (freckles):

so apply the product rule to find (xy)'

rvc (rvc):

yep :) you may have a look at that

rvc (rvc):

good then go for it apply the product rule

OpenStudy (lili.gallegos):

Okay thank you !

rvc (rvc):

All the best